\nocite{*}
\begin{thebibliography}{}



\bibitem{Bal}
\newblock
O. Balkanova, \emph{The first moment of Maass form symmetric square $L$-functions}, Ramanujan J., 55 (2021), 761--781.

\bibitem{BF2}
\newblock
O. Balkanova, D. Frolenkov, \emph{
Convolution formula for the sums of generalized Dirichlet L-functions}, Rev. Mat. Iberoam. 35:7 (2019), 1973--1995.

\bibitem{BF2mom}
\newblock
O. Balkanova,  D. Frolenkov, \emph{Non-vanishing of Maass form symmetric square $L$-functions}, J. Math. Anal. Appl., 500:2 (2021), 125148 , 23 pp.

%\bibitem{BFJEMS}
%\newblock
%O. Balkanova,  D. Frolenkov, \emph{Moments of L-Functions and the Liouville–Green Method},
%Journal of the European Mathematical Society 23:4 (2021), 1333–1380.


\bibitem{BFHanWu}
\newblock
O. Balkanova, D. Frolenkov, and H.Wu, \emph{On Weyl's subconvex bound for cube-free Hecke characters: totally real case}, arXiv:2108.12283, 2021.


\bibitem{BE}
\newblock
H. Beitman and A. Erdelyi, \emph{Higher transcendental functions}, Vol. 1, McGraw-Hill, New York, 1953.




\bibitem{BD}
\newblock
W.G.C. Boyd and T.M. Dunster, \emph{Uniform asymptotic solutions of a class of second-order linear differential equations having a turning point and a regular singularity, with an application to Legendre functions}, SIAM J. Math. Anal. 17-2 (1986), 422-450.



\bibitem{Dun90}
\newblock
T.M. Dunster, \emph{Uniform asymptotic solutions of second-order linear differential equations
having a double pole with complex exponent and a coalescing turning point}, SIAM J. Math.
Anal., 21 (6) (1990) 1594-1618.

%\bibitem{Dun90Bessel}
%\newblock
%T.M. Dunster, \emph{Bessel functions of purely imaginary order with
%applications to second order linear differential equations}, SIAM J. Math. Anal., 21 (4), (1990), pp.995-1018.


%\bibitem{Dun03}
%\newblock
%T.M. Dunster, \emph{Uniform asymptotic expansions for associated Legendre functions of large order}, Proc. Roy.
%Soc. Edinburgh Sect. A, 133 (2003), pp. 807–827,

%\bibitem{Dun13}
%\newblock
%T.M. Dunster, \emph{Conical functions of purely imaginary order and argument}, Proceedings of the Royal
%Society of Edinburgh 143A (2013), 929–955.


%\bibitem{KD2}
%\newblock
%S. Farid Khwaja and A. B. Olde Daalhuis, \emph{Uniform asymptotic expansions for hypergeometric functions with large parameters IV}, Anal. Appl. (Singap.), 12 (2014), pp. 667--710, https://doi.org/10.1142/S0219530514500389.


\bibitem{GR}
\newblock
I.S. Gradshteyn and I.M. Ryzhik, \emph{ Table of Integrals, Series, and
Products}. Edited by A. Jeffrey and D. Zwillinger. Academic Press, New York,
7th edition, 2007.



%\bibitem{Lam}
%\newblock
%J. W. C. Lam, \emph{The second moment of the central values of the symmetric square $L$-functions}, Ramanujan J. 38(1) (2015), 129–-145.


%\bibitem{Khan}
%\newblock
%R. Khan,  \emph{Non-vanishing of the symmetric square $L$-function at the central point.} Proc. London Math. Soc. (2010), 100 (3): 736--762.

%\bibitem{KhYoung}
%\newblock
%R. Khan, M. Young \emph{Moments and hybrid subconvexity for symmetric-square $L$-functions},
%Journal of the Institute of Mathematics of Jussieu (Dec. 2021), pp. 1–-45. DOI: 10.1017/S1474748021000566.




\bibitem{Olver}
\newblock
F.W.J. Olver, \emph{Asymptotics and Special Functions}, Academic Press, New York, 1974.

%\bibitem{Olver75}
%\newblock
%F. W. J. Olver, \emph{Legendre functions with both parameters large}, Philos. Trans. Roy. Soc. London Ser. A, 278
%(1975), pp. 175–185

\bibitem{HMF}
\newblock
F.W.J. Olver , D.W. Lozier, R.F. Boisvert and C.W. Clarke, \emph{{NIST}
{H}andbook of {M}athematical {F}unctions}, Cambridge University Press,
Cambridge 2010.


%\bibitem{TangXu}
%\newblock
%H. Tang, Z. Xu, \emph{Central value of the symmetric square $L$-functions related to Hecke-Maass forms} Lith. Math. J. 56.2 (2016), 251--267.


\bibitem{Temme1983}
N. M. Temme, \emph{ Uniform  asymptotic  expansions  of Laplace  integrals}, Analysis, 3,  221-249

\bibitem{Temme1985}
N. M. Temme, \emph{Laplace type intergals: Transformation to standard form and uniform asymptotic expansions}. Quart. Appl. Math. 43(1):103123, 1985.

\bibitem{Temme2015}
N. M. Temme, \emph{Asymptotic methods for integrals}, volume 6 of Series in Analysis. World
Scientific Publishing Co. Pte. Ltd., Hackensack, 2015.

\bibitem{Temme2021}
N. M. Temme, \emph{Asymptotic expansions of Kummer hypergeometric functions for large values of parameters}, Integral Transforms and Special Functions 33.1 (2022): 16-31.




\bibitem{Zav}
\newblock
N. I. Zavorotny, \emph{On the fourth moment of the Riemann zeta-function}, Automorphic functions and number theory
part I, II (Russian), Akad. Nauk SSSR, Dalnevostochn. Otdel., Vladivostok 254 (1989), p. 69--124.


\end{thebibliography}

\documentclass[12pt]{amsart}
\usepackage{amssymb}
\usepackage{verbatim}
\usepackage{mathrsfs}
\usepackage{mathtools}

\usepackage{xcolor}

\usepackage{caption}

\usepackage{blindtext}

\newtheorem{thm}{Theorem}[section]

\theoremstyle{definition}

\theoremstyle{remark}

\numberwithin{equation}{section}

\providecommand{\arcosh}{\operatorname{arcosh}}
\newcommand{\alt}{\alpha}
\newcommand{\Lo}{ l_1 }

\newcommand{\HyG}{ {}_2F_1 }

\mathtoolsset{showonlyrefs}

\begin{document}

\title[Asymptotics of a Gauss hypergeometric function ]{Asymptotics of a Gauss hypergeometric function related to moments of symmetric square $L$-functions I}

\author[O.Balkanova]{Olga Balkanova}
\address{
Steklov Mathematical Institute of Russian Academy of Sciences, 8 Gubkina st., Moscow, 119991, Russia}
\email{balkanova@mi-ras.ru}


\begin{abstract}
We prove an asymptotic formula for a special case of the Gauss hypergeometric function which arises in explicit formulas for moments of Maass form symmetric square $L$-functions. The resulting formula is uniform in several variables, which is crucial for proving hybrid subconvexity bounds for the $L$-functions under consideration.
\end{abstract}

\keywords{Gauss hypergeometric function, saddle point method}
\subjclass[2010]{Primary:33C05, 41A60}

\maketitle


\section{Introduction}
This paper is devoted to proving an asymptotic formula for the function
\begin{equation}\label{2f1(1/4 3/4 1)}
\HyG\left(1/4+ir(1-\alt),3/4+ir(1-\alt),1+2ir;z\right),\quad 0<z<1,\,
r\to\infty, \,\alpha=t/r\to0.
\end{equation}
This is one of the two special cases of the Gauss hypergeometric function appearing in the explicit formula for  the first moment of Maass form symmetric square $L$-functions on the critical line $\rho=1/2+2it$, see \cite[Theorem 5, Lemma 7]{Bal} for details. Our goal is to use this explicit formula as a starting point to study the second moment on short intervals, and ultimately to prove new hybrid subconvexity bounds for symmetric square $L$-functions. To this end, it is required to have the following uniform expansion for \eqref{2f1(1/4 3/4 1)}.

\begin{thm}\label{prop: 2F1 asympt1}
Assume that $0<z<1,$ $r\to\infty$ and $r^{-1+\delta}<\alt< r^{-\delta}$. Then 
\begin{multline}\label{2F1 main asymptI}
\HyG\left(1/4+ir(1-\alt),3/4+ir(1-\alt),1+2ir;z\right)=\\=
\frac{\exp\left(2ir\Lo(\alt,z)\right)}{(1-(1-\alt^2)z)^{1/4}}
\left(1+\sum_{j=1}^{N}\frac{c_j(\alt,r)}{(\alt r)^j}\right)+O((\alt r)^{-N-1}),
\end{multline}
where
\begin{equation}\label{2F1 main asymptI Lo func}
\Lo(\alt,z)=\log2-\alt\log(1+\alt)-\log(1+\sqrt{1-(1-\alt^2)z})+\alt\log(\alt+\sqrt{1-(1-\alt^2)z}).
\end{equation}
\end{thm}
This asymptotic formula generalizes \cite[Lemma 4.6]{BF2mom},  where the case $\alt=0$ has been considered. Note that for $\alt=0$ and $z=4/x^2$ the main term in \eqref{2F1 main asymptI}
reduces to the expression (see \cite[4.37.19]{HMF})
\begin{equation}
\frac{\exp\left(2ir\left(\log2-\log(1+\sqrt{1-4/x^2})\right)\right)}{(1-4/x^2)^{1/4}}=
x^{2ir}\exp(-2ir\arcosh{x/2})
\left(\frac{x^2}{x^2-4}\right)^{1/4},
\end{equation}
which coincides with the main term in \cite[Lemma 4.6]{BF2mom}.
\section{The first case: $0<z<1-\delta$}
We start the proof of \eqref{2F1 main asymptI} by considering the case $0<z<1-\delta$.  In these settings, several methods may be used to derive the required asymptotic formula, including the Liouville-Green method  (by generalizing \cite[Lemma 2.4]{Zav} and \cite[Corollary 7.3]{BF2}) or a standard  version of the saddle point method (generalizing \cite[Lemma 20]{Bal}). However, the common problem of these two methods is that they does not seem to be applicable when $z$ is very close to $1$. Let us provide some explanation related to this issue starting with the Liouville-Green method and later discussing the saddle point method.

The Liouville-Green method is based on studying solutions of differential equations, see \cite[Ch. 6, pp. 10--12]{Olver}.   More precisely, using \cite[p.96, f.(7)-(9)]{BE} we conclude that
the function
\begin{equation}\label{Y def}
Y(z)=z^{1/2+ir}(1-z)^{1/2-ir\alt}\HyG\left(1/4+ir(1-\alt),3/4+ir(1-\alt),1+2ir;z\right)
\end{equation}
satisfies the differential equation
\begin{equation}\label{Y difeq}
Y''(z)=\left((ir)^2h(\alt,z)+g(\alt,z)\right)Y(z),
\end{equation}
where
\begin{equation}
h(\alt,z)=\frac{1-(1-\alt^2)z}{z^2(1-z)^2},\,
g(\alt,z)=-\frac{1-3z/4+3z^2/4}{4z^2(1-z)^2}.
\end{equation}
The  function $h(\alt,z)$ has double poles at the points $0$ and $1$ and a turning point at $z_0=(1-\alt^2)^{-1}$. If $z$ is not to close to $1$, then  this turning point does not play any role, and consequently, it is possible to apply the theory developed in  \cite[Ch.6 Sec. 5.3, Ch.10 Sec. 4.1]{Olver}.  According to \cite[Ch.10, Eqs. (1.01), (2.01)-(2.03), (2.09), (2.10)]{Olver}
and \cite[Ch.10, Thms. 3.1 and 4.1]{Olver} one has
\begin{equation}\label{Y to W1W2}
Y(z)h^{1/4}(\alt,z)=C_1W_{n,1}(r,\xi)+C_2W_{n,2}(r,\xi),
\end{equation}
where
\begin{equation}\label{xi def}
\xi=\int\sqrt{h(\alt,z)}dz,
\end{equation}
and
\begin{equation}\label{W1W2 def}
W_{n,1}(r,\xi)=e^{ir\xi}\sum_{j=0}^{n-1}\frac{A_{j}(z)}{(ir)^j}+O(r^{-n}),\quad
W_{n,2}(r,\xi)=e^{-ir\xi}\sum_{j=0}^{n-1}\frac{(-1)^jA_{j}(z)}{(ir)^j}+O(r^{-n}),
\end{equation}
\begin{equation}\label{Aj def}
A_{j+1}(z)=\frac{-A'_j(z)}{2\sqrt{h(\alt,z)}}+\int\Lambda(z)A_j(z)dz,
\end{equation}
\begin{equation}\label{Lambda def}
\Lambda(z)=\frac{16g(\alt,z)h^2(\alt,z)+4h(\alt,z)h''(\alt,z)-5(h'(\alt,z))^2}{32h^{5/2}(\alt,z)}, \quad A_0(z)=1.
\end{equation}
Applying  \cite[2.225.1, 2.224.5]{GR} we find that
\begin{multline}\label{xi eq}
\xi=-2\log(1+\sqrt{1-(1-\alt^2)z})+2\alt\log(\alt+\sqrt{1-(1-\alt^2)z})+\\+\log(z)-\alt\log(1-z)+
(1-\alt)\log(1-\alt^2).
\end{multline}
In order to determine $C_1$ and $C_2$ we compare two sides of equation \eqref{Y to W1W2} as $z\to0$. By \eqref{Y def} one has $Y(z)h^{1/4}(\alt,z)\sim z^{ir}$. As  a result, using  \eqref{xi eq} we infer that $C_2=0$ and
\begin{equation}\label{C1 def}
C_1=e^{ir\left(2\log2-2\alt\log(1+\alt)-(1-\alt)\log(1-\alt^2)\right)}.
\end{equation}
Finally,
\begin{multline}\label{2F1 LG asympt}
\HyG\left(1/4+ir(1-\alt),3/4+ir(1-\alt),1+2ir;z\right)=\\=
\frac{e^{2ir\Lo(\alt,z)}}{(1-(1-\alt^2)z)^{1/4}}
\left(1+\sum_{j=1}^{n-1}\frac{A_{j}(z)}{(ir)^j}\right)+O(r^{-n}),
\end{multline}
where $\Lo(\alt,z)$ is given by \eqref{2F1 main asymptI Lo func}, showing that \eqref{2F1 main asymptI} is proved for $z<1-\delta$.

The case when $z\to1$ is more complicated since in this situation the turning point coalesce with a pole as $\alpha\to0$. Asymptotic solutions of such differential equations have been investigated by Boyd and Dunster \cite{BD} and Dunster \cite{Dun90}. In particular, the differential equation \eqref{Y difeq} satisfies the conditions in \cite{Dun90}.  Nevertheless,  it is not obvious how to apply the results of \cite{Dun90} in order to study \eqref{Y def}. The problem is that in \cite{Dun90} only real solutions of the differential equation have been considered and our function $Y(z)$ is not real. Note that Dunster \cite{Dun90} applied his result to the study of Ferrers functions, which are not real in general. For this reason,  instead of individual functions Dunster studied their tricky combination \cite[(5.15)]{Dun90}, which ultimately is a real function.

Being unable to overcome this difficulty in our case, we decided to use  the saddle point method for proving \eqref{2F1 main asymptI} for values of $z$ that are close to one. Let us start with the case $z<1-\delta$.
Applying \cite[15.6.1]{HMF} one has
\begin{multline}
\HyG\left(1/4+ir(1-\alt),3/4+ir(1-\alt),1+2ir;z\right)=\\=
\frac{\Gamma(1+2ir)}{\Gamma\left(3/4+ir(1-\alt)\right)\Gamma\left(1/4+ir(1+\alt)\right)}\int_0^1
\frac{y^{-1/4}\exp(irf(\alt,y))}{(1-y)^{3/4}(1-zy)^{1/4}}dy,
\end{multline}
where
\begin{equation}
f(\alt,y)=(1-\alt)\log y+(1+\alt)\log(1-y)-(1-\alt)\log(1-zy).
\end{equation}
The saddle points (solutions of $\frac{\partial}{\partial y}f(\alt,y)=0$) are given by
$$y_{\pm}=\frac{1\pm\sqrt{1-(1-\alt^2)z}}{(1+\alt)z}.$$
For $z\to1$ and $\alt\to0$ both $y_{+}$ and $y_{-}$ tend to the end point of the segment of integration and coalesce if $z=1,\alt=0$. This location of saddle points makes it impossible to apply the standard saddle point method.

For $0<z<1-\delta$ the point $y_{+}$ lies far away from the interval of integration, while the point $y_{-}$ is located inside the interval of integration and is not close to the end point.  In this case, the standard saddle point method yields the following asymptotic expansion:
\begin{multline}\label{2f1 sadpoint aprox}
\frac{\Gamma\left(3/4+ir(1-\alt)\right)\Gamma\left(1/4+ir(1+\alt)\right)}{\Gamma(1+2ir)}
\HyG\left(1/4+ir(1-\alt),3/4+ir(1-\alt),1+2ir;z\right)=\\=
\frac{y_{-}^{-1/4}e^{irf(\alt,y_{-})+\pi i/4}
\sqrt{2\pi}}{(1-y_{-})^{3/4}(1-zy_{-})^{1/4}\sqrt{rf''(\alt,y_{-})}}
\left(1+\sum_{j=1}^{N-1}\frac{c_j}{r^j}\right)+O(r^{-N-1/2}).
\end{multline}
Direct computations show that
\begin{multline}\label{f at y-}
f(\alt,y_{-})=(1-\alt)\log(1-\alt^2)-2\log(1+\sqrt{1-(1-\alt^2)z})+2\alt\log(\alt+\sqrt{1-(1-\alt^2)z}),
\end{multline}
\begin{equation}\label{g at y-}
\frac{y_{-}^{-1/4}(1-y_{-})^{-3/4}}{(1-zy_{-})^{1/4}}=\frac{(1+\alt)^{1/4}(1+\sqrt{1-(1-\alt^2)z})}{(1-\alt)^{1/4}(\alt+\sqrt{1-(1-\alt^2)z})},
\end{equation}
\begin{equation}\label{f'' at y-}
f''(\alt,y_{-})=-2\frac{(1+\alt)\left(1+\sqrt{1-(1-\alt^2)z}\right)^2\sqrt{1-(1-\alt^2)z}}{(1-\alt)\left(\alt+\sqrt{1-(1-\alt^2)z}\right)^2}.
\end{equation}
Finally, using the Stirling formula we prove that
\begin{multline}\label{Stirling2}
\Gamma(\sigma+it)=\sqrt{2\pi}|t|^{\sigma-1/2}\exp(-\pi|t|/2)
\exp\left(i\left(t\log|t|-t+\frac{\pi t(\sigma-1/2)}{2|t|}\right)\right)\\\times
\left(1+\sum_{j=1}^{N-1}a_j/t^j+O(|t|^{-N})\right),
\end{multline}
which holds for $|t|\rightarrow\infty$ and a  fixed $\sigma$. Note that
\begin{multline}\label{2f1 sadpoint Gamma}
\frac{\Gamma(1+2ir)}{\Gamma\left(3/4+ir(1-\alt)\right)\Gamma\left(1/4+ir(1+\alt)\right)}=O(r^{-N+1/2})+\\+
\frac{(1+\alt)^{1/4}\sqrt{r}}{(1-\alt)^{1/4}\sqrt{\pi}}
e^{\pi i/4+ir\left(2\log2-(1+\alt)\log(1+\alt)-(1-\alt)\log(1-\alt)\right)}
\left(1+\sum_{j=1}^{N-1}\frac{c_j}{r^j}\right).
\end{multline}
Substituting \eqref{f at y-}, \eqref{2f1 sadpoint Gamma} to \eqref{2f1 sadpoint aprox} we complete the proof of \eqref{2F1 main asymptI}.
\section{The second case: $1-\delta<z<1$}
It is left to prove \eqref{2F1 main asymptI} for $1-\delta<z<1$. First, we transform \eqref{2f1(1/4 3/4 1)} in the same way as in
\cite[Lemma 4.6]{BF2mom}. Namely, using \cite[Eq. 16, p. 112]{BE} and \cite[Eq. 15.8.1]{HMF}, we show that
\begin{multline}\label{2f1 double transform}
\HyG\left(1/4+ir(1-\alt),3/4+ir(1-\alt),1+2ir;z\right)\\=
(1-z)^{-1/4-ir(1-\alt)}\HyG\left( \frac{1}{2}+2ir(1-\alt),\frac{1}{2}+2ir(1+\alt),1+2ir; \frac{\sqrt{1-z}-1}{2\sqrt{1-z}}\right)\\
=(1-z)^{-1/4-ir(1-\alt)}\left(\frac{\sqrt{1-z}+1}{2\sqrt{1-z}}\right)^{-2ir}
\HyG\left( \frac{1}{2}+2ir\alt,\frac{1}{2}-2ir\alt,1+2ir; \frac{\sqrt{1-z}-1}{2\sqrt{1-z}}\right).
\end{multline}
Let $Y=\frac{1-\sqrt{1-z}}{2\sqrt{1-z}}$. Note that $Y\to+\infty$ as $z\to1$. Applying \cite[15.6.1]{HMF} we infer
\begin{multline}
\HyG\left( \frac{1}{2}+2ir\alt,\frac{1}{2}-2ir\alt,1+2ir; -Y\right)=\\=
\frac{\Gamma(1+2ir)}{\Gamma\left(1/2+2ir\alt\right)\Gamma\left(1/2+2ir(1-\alt)\right)}\int_0^1
\frac{t^{-1/2+2ir\alt}(1-t)^{-1/2+2ir(1-\alt)}}{(1+Yt)^{1/2-2ir\alt}}dt.
\end{multline}
Next, we make the change of variables: $t:=1-e^{-x}$. As a result,
\begin{multline}\label{2f1 to Temme int}
\HyG\left(1/4+ir(1-\alt),3/4+ir(1-\alt),1+2ir;z\right)=\\=
\frac{(1-z)^{-1/4-ir(1-\alt)}(1+Y)^{-2ir}\Gamma(1+2ir)}{\Gamma\left(1/2+2ir\alt\right)\Gamma\left(1/2+2ir(1-\alt)\right)}\int_0^{\infty}
e^{2ir(-x+\alt\log q(x))}\frac{dx}{\sqrt{q(x)}},
\end{multline}
where
\begin{equation}\label{q(x) def}
q(x)=\left(1-e^{-x}\right)\left((Y+1)e^x-Y\right).
\end{equation}
Let $w:=2ir, \lambda:=\alt w=2it$, then the integral on the right-hand side of \eqref{2f1 to Temme int} has the following form:
\begin{equation}\label{Temme int 1}
\int_0^{\infty}e^{-w(x-\alt\log q(x))}\frac{dx}{\sqrt{q(x)}}=
\int_0^{\infty}q(x)^{\lambda-1}e^{-wx}\sqrt{q(x)}dx.
\end{equation}
Such integrals were studied  by  Temme  \cite{Temme1983}, \cite{Temme1985}, \cite[Ch. 25]{Temme2015}, \cite{Temme2021}, and also appeared in \cite{BFHanWu}.  The saddle point of the integral \eqref{Temme int 1} (solution of $q(x)=\alt q'(x)$) is given by
\begin{equation}\label{Temme x0 def}
x_0=\log\frac{1+\sqrt{1-(1-\alt^2)z}}{(1-\alt)\left(1+\sqrt{1-z}\right)}.
\end{equation}
We perform all further transformations assuming that $\Re{w},\Re{\lambda}>0$, and then apply the technique of analytic continuation to obtain the required result in the region of interest.  The first step of Temme's method is to make the change of variables:
\begin{equation}\label{Temme change of variable}
x-\alt\log q(x)=t-\alt\log t+A(\alt).
\end{equation}
This transformation should move the point $x=0$ to $t=0,$ the point $x=x_0$ to $t=\alt$, and the point $x=+\infty$ to $t=+\infty$. The regularity and other properties of this transformation are proved in \cite[Sec. 2]{Temme1985}. We remark (see \cite[(2.5), (2.7), (2.15)]{Temme1985}) that
\begin{equation}\label{Temme dx/dt}
\frac{dx}{dt}=\frac{q(x)(t-\alt)}{t\left(q(x)-\alt q'(x)\right)},
\end{equation}
\begin{equation}\label{Temme dt/dx}
\frac{dt}{dx}\Biggl|_{x_0}=\sqrt{1-\alt^2\frac{q''(x_0)}{q(x_0)}},
\end{equation}
\begin{equation}\label{Temme Adef}
A(\alt)=x_0-\alt\log q(x_0)-\alt+\alt\log\alt.
\end{equation}
After performing the change of variable one has
\begin{equation}\label{Temme int 2}
\int_0^{\infty}e^{-w(x-\alt\log q(x))}\frac{dx}{\sqrt{q(x)}}=
e^{-wA(\alt)}\int_0^{\infty}e^{-w(t-\alt\log t)}f(t)\frac{dt}{t},\quad
f(t)=\frac{t}{\sqrt{q(x)}}\frac{dx}{dt}.
\end{equation}
The next step of Temme's method \cite[Sec. 5]{Temme1983}, which is a modification of the procedure developed by Franklin and Friedman \cite[Sec. 4]{Temme1983}, is to write
\begin{equation}\label{Temme f transform}
f(t)=f(\alt)+(f(t)-f(\alt)),
\end{equation}
and to perform integration by parts, resulting in
\begin{equation}\label{Temme int 3}
\int_0^{\infty}e^{-w(t-\alt\log t)}f(t)\frac{dt}{t}=
f(\alt)\frac{\Gamma(\lambda)}{w^{\lambda}}+\frac{1}{z}\int_0^{\infty}e^{-w(t-\alt\log t)}\widetilde{f}_1(t)\frac{dt}{t},
\end{equation}
where
\begin{equation}\label{Temme fi tilde}
\widetilde{f}_1(t)=t\left(\frac{f(t)-f(\alt)}{t-\alt}\right)'.
\end{equation}
Repeating this process we show that
\begin{equation}\label{Temme int 3}
\int_0^{\infty}e^{-w(t-\alt\log t)}f(t)\frac{dt}{t}=
\Gamma(\lambda)\sum_{k=0}^{n-1}\frac{\widetilde{f}_k(\alt)}{w^{\lambda+k}}+\frac{1}{w^n}\int_0^{\infty}e^{-w(t-\alt\log t)}\widetilde{f}_n(t)\frac{dt}{t},
\end{equation}
where
\begin{equation}\label{Temme fk tilde}
\widetilde{f}_0(t)=f(t),\quad \widetilde{f}_{k+1}(t)=t\left(\frac{\widetilde{f}_k(t)-\widetilde{f}_k(\alt)}{t-\alt}\right)'.
\end{equation}
Note that the values $\widetilde{f}_k(\alt)$ can be evaluated in terms of  Taylor series coefficients for the function $f(t)$:
\begin{equation}\label{Temme f Taylor}
f(t)=\sum_{j=0}^{\infty}a_j(\alt)(t-\alt)^j.
\end{equation}
According to  \cite[(3.35)]{Temme1985} and  \cite[(25.1.6), (25.2.26)]{Temme2015}
\begin{equation}\label{Temme fk tilde(alt)}
\widetilde{f}_0(\alt)=a_0(\alt),\quad
\widetilde{f}_1(\alt)=\alt a_2(\alt),\quad
\widetilde{f}_2(\alt)=\alt\left(3\alt a_4(\alt)+2a_3(\alt)\right).
\end{equation}

To estimate the integral in \eqref{Temme int 3} we argue in the same way as in \cite[Lemma 6.18]{BFHanWu}. First, we split the integral into two parts:
$|t-\alt|<\alt/100$ and $|t-\alt|>\alt/100$. When $t$ is close to $\alt$ we use \cite[pp. 241-242]{Temme1983}:
\begin{equation}\label{Temme fk to f k-1}
\widetilde{f}_k(t)=\frac{t}{2}\widetilde{f}_{k-1}''(\xi),\quad \hbox{where}\quad\xi\quad\hbox{lies between}\quad t\quad\hbox{and}\quad\alt,
\end{equation}
\begin{equation}\label{Temme fk tilde near alt}
\widetilde{f}_k(t)=t\sum_{r=1}^{k}a_k^{(r)}\alt^{r-1}f^{(r+k)}(t_r),\quad \hbox{where}\quad t_r\quad\hbox{lies between}\quad t\quad\hbox{and}\quad \alt,
\end{equation}
where $a_k^{(r)}\ll1.$ Since $f(t)\sim t^{1/2}$ for small $t$ we conclude that  $\widetilde{f}_k(t)\ll \alt^{1/2-k}$ if $t$ close to $\alt$. Therefore,
\begin{equation}\label{Temme int est1}
\frac{1}{w^n}\int_{|t-\alt|<\alt/100}e^{-w(t-\alt\log t)}\widetilde{f}_n(t)\frac{dt}{t}\ll\frac{1}{(w\alt)^{n-1}w\sqrt{\alt}}.
\end{equation}
If $t\to0$ then arguing inductively based on \eqref{Temme fk tilde} one has $\widetilde{f}_k(t)\ll\alt^{-k}\sqrt{t}$. As a result,  since $\widetilde{f}_n(t)$ decays exponentially at infinity the following estimate holds:
\begin{equation}\label{Temme int est2}
\frac{1}{w^n}\int_{|t-\alt|>\alt/100}e^{-w(t-\alt\log t)}\widetilde{f}_n(t)\frac{dt}{t}\ll\frac{1}{(w\alt)^{n-1}w\sqrt{\alt}}.
\end{equation}
Using  \eqref{Temme int 3}, \eqref{Temme int est1} and \eqref{Temme int est2}  we obtain
\begin{equation}\label{Temme int 4}
\int_0^{\infty}e^{-w(t-\alt\log t)}f(t)\frac{dt}{t}=
\Gamma(\lambda)\sum_{k=0}^{n-1}\frac{\widetilde{f}_k(\alt)}{w^{\lambda+k}}+O\left(\frac{1}{(w\alt)^{n-1}w\sqrt{\alt}}\right).
\end{equation}
Applying \eqref{Temme int 4} to the integral in \eqref{2f1 to Temme int} (recall that $w=2ir, \lambda=\alt w=2it$) one has
\begin{equation}\label{2f1 to Temme int2}
\int_0^{\infty}e^{2ir(-x+\alt\log q(x))}\frac{dx}{\sqrt{q(x)}}=
e^{-2irA(\alt)}\frac{\Gamma(2it)}{(2ir)^{2it}}\sum_{k=0}^{n-1}\frac{\widetilde{f}_k(\alt)}{(2ir)^{k}}+O\left(\frac{1}{(r\alt)^{n-1}r\sqrt{\alt}}\right).
\end{equation}
It is left to evaluate the main term and to prove that it coincides with the one in \eqref{2F1 main asymptI}. According to \eqref{Temme dt/dx}, \eqref{Temme Adef} and \eqref{Temme int 2}  the main term in \eqref{2f1 to Temme int2} is equal to
\begin{equation}\label{Temme mainterm0}
e^{-2ir\left(x_0-\alt\log q(x_0)-\alt+\alt\log\alt\right)}\frac{\Gamma(2it)}{(2ir)^{2it}}\frac{\alt}{\sqrt{q(x_0)-\alt^2q''(x_0)}},
\end{equation}
where $x_0$ is defined by \eqref{Temme x0 def}, and $q(x)$ is given by \eqref{q(x) def}. Using \eqref{Stirling2} one has
\begin{equation}\label{Temme mainterm1}
\frac{\Gamma(2it)}{(2ir)^{2it}}=\frac{\sqrt{\pi}}{\sqrt{r\alt}}e^{-\pi i/4+2ir(\alt\log\alt-\alt)}\left(1+\sum_{j=1}^{N-1}\frac{c_j}{t^j}\right)+O(t^{-N-1/2}).
\end{equation}
Straightforward computations show that
\begin{equation}\label{Temme mainterm2}
q(x_0)-\alt^2q''(x_0)=\frac{\alt\sqrt{1-(1-\alt^2)z}}{\sqrt{1-z}},
\end{equation}
\begin{equation}\label{Temme mainterm3}
x_0-\alt\log q(x_0)-\alt+\alt\log\alt=\log\frac{1+\sqrt{1-(1-\alt^2)z}}{(1-\alt)(1+\sqrt{1-z})}-
\alt\log\frac{\alt+\sqrt{1-(1-\alt^2)z}}{(1-\alt^2)\sqrt{1-z}}-\alt.
\end{equation}
Substituting \eqref{Temme mainterm1}, \eqref{Temme mainterm2} and \eqref{Temme mainterm3} to \eqref{Temme mainterm0}  we conclude that the main term in \eqref{2f1 to Temme int2} is given by
\begin{equation}\label{2f1 to Temme int3}
\frac{e^{-\pi i/4}\sqrt{\pi}}{\sqrt{r}}\frac{(1-z)^{1/4}}{(1-(1-\alt^2)z)^{1/4}}e^{2ir\left(
\alt\log\alt-\log\frac{1+\sqrt{1-(1-\alt^2)z}}{(1-\alt)(1+\sqrt{1-z})}+
\alt\log\frac{\alt+\sqrt{1-(1-\alt^2)z}}{(1-\alt^2)\sqrt{1-z}}
\right)}.
\end{equation}
Substituting  \eqref{2f1 to Temme int2} and \eqref{2f1 to Temme int3} to  \eqref{2f1 to Temme int}, applying
\begin{equation}
\frac{\Gamma(1+2ir)}{\Gamma\left(1/2+2ir\alt\right)\Gamma\left(1/2+2ir(1-\alt)\right)}
=\frac{\sqrt{r}}{\sqrt{\pi}}e^{\pi i/4+2ir(-\alt\log\alt-(1-\alt)\log(1-\alt))}\left(1+\sum_{j=1}^{N-1}\frac{c_j}{t^j}\right)+O(t^{-N-1/2}),
\end{equation}
and using the fact that $Y=\frac{1-\sqrt{1-z}}{2\sqrt{1-z}}$, we obtain
\begin{multline}\label{2f1 Temme asympt1}
\HyG\left(1/4+ir(1-\alt),3/4+ir(1-\alt),1+2ir;z\right)=\\=
\frac{e^{2irF(\alt,z)}}{(1-(1-\alt^2)z)^{1/4}}\left(1+\sum_{j=1}^{N-1}\frac{c_j}{t^j}\right)+O(t^{-N}),
\end{multline}
where
\begin{multline}\label{2f1 to Temme mt F def}
F(\alt,z)=-\log\frac{1+\sqrt{1-(1-\alt^2)z}}{(1-\alt)(1+\sqrt{1-z})}+
\alt\log\frac{\alt+\sqrt{1-(1-\alt^2)z}}{(1-\alt^2)\sqrt{1-z}}-\\-
\log\frac{1+\sqrt{1-z}}{2\sqrt{1-z}}-(1-\alt)\log\sqrt{1-z}-(1-\alt)\log(1-\alt).
\end{multline}
One can easily check that $F(\alt,z)$ coincides with $\Lo(\alt,z)$  defined in \eqref{2F1 main asymptI Lo func}. This completes the proof of \eqref{2F1 main asymptI}.


\section{Numerical examples}

In this section, we provide some numerical examples related to Theorem \ref{prop: 2F1 asympt1}.
More precisely, we compute for some values of $r,$ $\alpha$ and $z$ the left-hand side of \eqref{2F1 main asymptI}, given by
$$F(r,\alpha,z):=\HyG\left(1/4+ir(1-\alt),3/4+ir(1-\alt),1+2ir;z\right),$$ and compare it
with the first summand on the right-hand side of \eqref{2F1 main asymptI}:
$$R(r,\alpha,z):=\frac{\exp\left(2ir\Lo(\alt,z)\right)}{(1-(1-\alt^2)z)^{1/4}}.$$

The results are given in the Tables below.

\begin{center}
 \begin{tabular}{ |c|c|c| }
\hline 
Function & Approximation & $|$ Relative error $|$ \\ \hline
 &  &  \\  
$F(r,\alpha,z)$  & $2.611880247+0.5174226366i$ &  \\ 
&  &  \\  
$R(r,\alpha,z)$ & $2.611802650+0.5170096337i$ & $0.000157843$  \\ 
&  &  \\ \hline
\end{tabular}
\captionof{table}{The case $r=100$, $\alpha=0.1$, $y=0.99$.}\label{case4}
\end{center}


\begin{center}
 \begin{tabular}{ |c|c|c| }
\hline 
Function & Approximation & $|$ Relative error $|$ \\ \hline
 &  &  \\  
$F(r,\alpha,z)$  & $-2.595771772-1.792471289i$ &  \\ 
&  &  \\  
$R(r,\alpha,z)$ & $-2.587392554-1.804512206i$ & $0.004650302$  \\ 
&  &  \\ \hline
\end{tabular}
\captionof{table}{The case $r=100$, $\alpha=0.1$, $y=0.9999$.}\label{case4}
\end{center}


\begin{center}
 \begin{tabular}{ |c|c|c| }
\hline 
Function & Approximation & $|$ Relative error $|$ \\ \hline
 &  &  \\  
$F(r,\alpha,z)$  & $1.000002393+0.004050009913i$ &  \\ 
&  &  \\  
$R(r,\alpha,z)$ & $1.000002393+0.004050054037i$ & $7.576766037*10^{-9}$  \\ 
&  &  \\ \hline
\end{tabular}
\captionof{table}{The case $r=100$, $\alpha=0.1$, $y=0.00001$.}\label{case4}
\end{center}


\begin{center}
 \begin{tabular}{ |c|c|c| }
\hline 
Function & Approximation & $|$ Relative error $|$ \\ \hline
 &  &  \\  
$F(r,\alpha,z)$  & $-2.258587650-2.168919043i$ &  \\ 
&  &  \\  
$R(r,\alpha,z)$ & $-2.269908138-2.1575933584i$ & $0.005113882$  \\ 
&  &  \\ \hline
\end{tabular}
\captionof{table}{The case $r=100$, $\alpha=0.02$, $y=0.99$.}\label{case4}
\end{center}


\begin{center}
 \begin{tabular}{ |c|c|c| }
\hline 
Function & Approximation & $|$ Relative error $|$ \\ \hline
 &  &  \\  
$F(r,\alpha,z)$  & $-3.328488374+5.815264147i$ &  \\ 
&  &  \\  
$R(r,\alpha,z)$ & $-3.380721684+5.770084005i$ & $0.010307062$  \\ 
&  &  \\ \hline
\end{tabular}
\captionof{table}{The case $r=100$, $\alpha=0.02$, $y=0.9999$.}\label{case4}
\end{center}


\end{document}